The standard deviation for the sample numbers 8,9,10? I got 0.81 and rounded to 1
Answer by edjones(8007)

The mean is 9.
(8-9)^2+(9-9)^2+(10-9)^2
=1+0+1=2
2/n-1
=2/3-1
=2/2
=1 Variance
sqrt(1)=1 standard deviation
.
Ed